# Multi-modal Retrieval Augmented Multi-modal Generation: Datasets, Evaluation Metrics and Strong Baselines

### v3

@@ 4.4 Training Dataset Construction
As another contribution to the research community, we construct a high-quality training dataset to enhance the performance of small-scale but efficient 7B-8B LLMs and MLLMs on the M2RAG task. Specifically, we first utilize the state-of-the-art GPT-4o with multi-stage strategy to construct 3K triplets of ($Q$, $K_{\text{In-Doc}}$, $r$). [^5] Experimental results in Section 5.2 suggest this setup leads to the best performance. Then, our proposed multi-modal metrics are used to filter the high-quality samples. We do not utilize the text-modal metrics, since the time cost of Context Precision and Faithfulness metrics on 3K long generations are huge [^6] Please refer to Section Limitations for more explanation., and we plan to improve the efficiency of text-modal evaluation metrics and update the training data in future work. Finally, the dataset consists of 1.6K instances. More implementation details and statistical information of our training dataset could be found in Appendix D.
@@ 5.1 Validate Reliability of Evaluation Metrics
Image Coherence MLLM 0.56 0.36 0.48
@@
Image Helpfulness MLLM 0.67 0.48 0.57
@@
Image Reference MLLM 0.61 0.43 0.53
@@
Table 1: The Spearman’s correlation between MLLMs and human annotators. Max, Min and Avg. represents the maximum, minimum and average correlations with all human annotators except for themselves.
Since the reliability of text-modal metrics have been proven in RAG scenarios Es et al. (2023), we mainly focus on validating the reliability of our proposed MLLM-based multi-modal metrics: (1) Image Coherence; (2) Image Helpfulness; and (3) Image Reference. In practice, we randomly select 200 samples from the benchmark labeled by each metric, and three annotators independently scored these samples using the same scoring criteria as the model. Specifically, we calculate the Spearman correlation coefficients between two scorers, each can be human annotators or MLLMs. Table 1 reveals that MLLMs achieve comparable performance to inner-correlation of human annotators. GPT-4o is even better than human average on helpfulness metric (0.57 > 0.53). These results indicate that our proposed multi-modal evaluation metrics are good proxy for human annotators.
@@ 5.2 Overall Experimental Results
Reasoner Single OpenAI o3-mini 80.8 89.5 91.9 90.0 73.3 63.2 64.6 79.1 79.1
@@
We conduct comprehensive experiments on our benchmark with 12 representative models, including OpenAI o3-mini, GPT-4o, DeepSeek-V3, Step-1o, Llama-3.1 & Llama-3.2-Vision series, Qwen2.5 & Qwen2-VL series (see Appendix F for details). Due to the high evaluation cost, we sample 200 queries from the benchmark dataset by category to compare the performance of all models. The results are shown in Table 2 , and we can observe several findings.
@@
Table 2 indicates that the most capable open-source and closed-source LLMs for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o, using a multi-stage strategy with LLM backbones. Among MLLM backbones, GPT-4o and Step-1o outperform all open-source models. For example, Step-1o surpassing Qwen2-VL-72B-Instruct by a large margin (83.0 > 79.9). The reasoner OpenAI o3-mini significantly outperforms other models in the single-stage approach. Despite these strong performances, there remains significant room for improvement, as even the best-performing model falls short of the ideal score (83.8 < 100.0).
@@
Regarding modeling approaches, separate modeling Zhu et al. (2024) proves significantly weaker than joint modeling, with even the worst-performing joint approach (GPT-4o, 74.9) surpassing separate modeling (74.1), highlighting the importance of modeling multi-modal interactions.
@@
Notably, while being constrained by evaluation costs, to investigate whether these findings remain valid as the test data scale increases, we selected several representative models and evaluated them on all 750 queries. The results, presented in Appendix I, confirm that the findings remain valid.
@@ 5.3 Improvements from Fine-tuning
We further conduct the supervised fine-tuning to improve the 7B-8B LLMs and MLLMs capabilities in M2RAG task. Please refer to Appendix H for more fine-tuning implementation details. The evaluation results are illustrated in Table 2. It can be observed that our fine-tuned LLMs and MLLMs exhibit significant improvements on all multi-modal evaluation metrics and the overall score. For example, the average improvements on the overall score is 19.2%. Since the text-modal metrics are not utilized for constructing training dataset, our fine-tuned models underperforms the baselines on some textual evaluation metrics, particularly Fluency.
@@ 6.1 Does Topic Affect Performance?
Figure 3: Average overall score across 10 topics.
@@ 6.2 Is Description Quality Important?
As described in Section 4.3, to provide the LLMs with rich and accurate image content, each image is transcribed into a detailed textual description, with consideration of its original contextual information. In this subsection, we conduct ablation studies to investigate the contributions of these information. Table 3 reveals that model’s performance significantly declines by removing contextual information, particularly in the text-modal and multi-modal precision metrics. The absence of details in descriptions leads to a notable drop in Image Recall, highlighting the critical role of details in guiding image selection of LLMs. Overall, incorporating context is essential for achieving optimal image retrieval, while including detailed descriptions further enhances performance to a measurable extent.
@@
Table 3: Comparison of GPT-4o experiment and ablation study results. Text Avg. and Multi-Modal Precision Avg. represent the average values of four text-modal metrics and three multi-modal precision metrics (except for Recall), respectively. Recall and Overall scores for w/o Aux. Image is not reported due to changes in the images within the knowledge base.
@@ 6.3 Are Auxiliary Images Helpful?
Auxiliary images are collected for each query to address cases where highly relevant images are absent from the retrieved web pages. As illustrated in last row of Table 3 (w/o Aux. Image), removing these auxiliary images leads to a significant decline in generative performance in multi-modal metrics. This observation reflects the challenges in effectively integrating auxiliary images that lack contextual information with the textual content. Overall, the inclusion of auxiliary images significantly enriches the visual content and enhances the quality of the generated document.
@@ 7 Conclusion
The proposed evaluation metrics might fail to capture all potential aspects of M2RAG evaluation. For example, pairwise ranking is more appropriate for assessing overall performance than Likert scale in M2RAG task, which lacks a standard optimal answer. Besides, due to the requirement of using LLMs and MLLMs to evaluate the quality of long-text multi-modal content across 8 dimensions [^7] Please refer to Table 6 for more details about evaluation cost., this leads to huge evaluation costs in large-scale evaluation. To balance the affordability and reliability of experiments on our proposed benchmark, we introduce a quantity-first approach with 750 evaluation samples in this paper. In future work, we plan to distill and build efficient text-modal and multi-modal metrics to support larger-scale benchmark testing.
Appendix A Ethical Considerations
Most of the task inputs in our benchmark and training dataset are sourced from publicly available datasets, ensuring that they pose no harm to individuals or groups. Furthermore, the text generated by large language models (LLMs) and multi-modal language models (MLLMs) is carefully curated and processed by human annotators to safeguard privacy and confidentiality. No personally identifiable information (PII) is included. However, it is important to note that texts from the ELI5 dataset Fan et al. (2019) and multi-modal documents retrieved via Google Custom Search may contain harmful content or hate speech. Despite these potential risks, it is crucial to disclose the full scope of this research, as materials from ELI5 and Google Custom Search have been extensively used in safety research within the community. All annotators are compensated fairly, with an hourly wage of approximately \$4.76 USD, which exceeds the average hourly wage of \$3.13 USD on Amazon Mechanical Turk Hara et al. (2018).
Appendix B Query Collection
To simulate the real-world user query solving problem, we collect diverse and high-quality user queries from the ELI5 dataset Fan et al. (2019). The ELI5 corpus is particularly suited for our task due to its comprehensive collection of long-form, open-ended questions that necessitate detailed, multi-sentence responses. The diversity of topics in ELI5 presents an opportunity to reflect the capabilities of language models in the real-world scenarios. Subsequently, we conduct two steps to collect diverse and high-quality queries that are suitable for using multi-modal information as responses: (1) Query Filtering; and (2) Query Classification.
@@ 1. Society & Culture
Example: What is the significance of the Day of the Dead celebration in Mexican culture? 2. Science & Mathematics: Questions in this category often deal with complex concepts that can be enhanced with diagrams, charts, or images. These visual aids can help in understanding scientific phenomena or mathematical theories. Example: How does photosynthesis work in plants? 3. Health: Health-related questions often benefit from the use of images like anatomical diagrams, charts, or photos for better understanding of medical conditions, treatments, or fitness exercises. Example: What are the symptoms and treatment options for carpal tunnel syndrome? 4. Education & Reference: This category includes academic topics and general knowledge where explanatory images, instructional graphics, and reference tables can provide clarity and enhance learning. Example: What are the main parts of a plant cell and their functions? 5. Computers & Internet: Questions about technology, software, and internet usage can be more effectively answered with screenshots, infographics, and step-by-step visual guides. Example: How do I set up a Wi-Fi network at home? 6. Sports: Sports-related queries can be explained better with visual demonstrations, play diagrams, athlete photos, and equipment images to illustrate techniques, rules, or historical moments. Example: What are the basic rules of soccer? 7. Business & Finance: Financial and business concepts often involve data, charts, and visualizations that can make complex information more digestible and insightful. Example: How does compound interest work and what impact can it have on investments? 8. Entertainment & Music: This category includes questions about movies, TV shows, music, and celebrities where images of album covers, scene snapshots, or charts of music theory can enrich the content. Example: What was the cultural impact of the TV show ’Friends’?
@@ 11 Others 451
Table 4: Topics and volumes of the queries.
Appendix C Retrieval of Collected Elements
@@ 2. Text from a Web Page # Evaluation Guidelines
markdown
{webpage_piece} ‘‘‘ # Precautions
@@
Appendix D Dataset Statistics
D.1 Basic Statistics
The statistical information about our benchmark are shown in Table 5.
Item Range Value
Query benchmark 750
training 1605
Web Page per query 9.8
Image 12.0
Aux. Image 3.0
Text Element per web page 25.7
Image 1.2
Avg. Element Score text 1.5
image 3.1
aux. image 3.8
Table 5: Statistical information of our benchmark.
D.2 Data Distribution
@@
Appendix E Details of Evaluation Metrics
E.1 Prompt Templates
@@
markdown
{text} ‘‘‘ # Scoring Criteria
@@
markdown
{text} ‘‘‘ # Scoring Criteria
@@
markdown
{context_above} ‘‘‘ # Image
@@
markdown
{context_below} ‘‘‘ # Scoring Criteria and Examples
@@
markdown
{context_above} ‘‘‘ # Image
@@
markdown
{context_below} ‘‘‘ # Scoring Criteria and Examples
@@
markdown
{context_above} ‘‘‘ # Image
@@
markdown
{context_below} ‘‘‘ # Scoring Criteria
@@ E.2 Evaluation Costs
Appendix F Experimental Setup
@@
We evaluate the capabilities of some advanced open-source and closed-source LLMs and MLLMs for M2RAG task. The involved models are listed in Table 7. Besides, we also implement a separate approach, i.e., the MuRAR method with GPT-4o as the backbone model Zhu et al. (2024).
@@
For the generation process, we deployed open-source LLMs and MLLMs using vLLM Kwon et al. (2023) on Nvidia A100-SXM4-80GB GPUs. Models with 7B-11B parameters operate on a single GPU, while those exceeding 70B parameters are distributed across 4 GPUs using tensor parallelism. The context length is configured at 64k for the Llama series and 32k for the Qwen series. During generation, we use a top-$k$ selection method with $k$ set to 20 for filtering text elements from each webpage. The maximum number of auxiliary and web page specific images is restricted to 5, with a total input image cap at 10. Images are resized to 512 $\times$ 512 thumbnails for MLLMs processing. For evaluation, we implemented our customized metrics using the GPT-4o model to ensure the robust comprehensive assessment. Besides, the RAG metrics are implemented with RAGAS Es et al. (2023) by using the GPT-4o-mini model.
Appendix G Case Study

### v4

@@ 4.4 Training Dataset Construction
As another contribution to the research community, we construct a high-quality training dataset to enhance the performance of small-scale but efficient 7B-8B LLMs and MLLMs on the M2RAG task. Specifically, we first utilize the state-of-the-art GPT-4o with multi-stage strategy to construct 3K triplets of ($Q$, $K_{\text{In-Doc}}$, $r$). [^5] Experimental results in Section 5.2 suggest this setup leads to the best performance. Then, our proposed multi-modal metrics are used to filter the high-quality samples. We do not utilize the text-modal metrics, since the time cost of Context Precision and Faithfulness metrics on 3K long generations are huge [^6] Please refer to Section A for more explanation., and we plan to improve the efficiency of text-modal evaluation metrics and update the training data in future work. Finally, the dataset consists of 1.6K instances. More implementation details and statistical information of our training dataset could be found in Appendix E.
@@ 5.1 Validate Reliability of Evaluation Metrics
Table 1: The Spearman’s correlation between MLLMs and human annotators. Max, Min and Avg. represents the maximum, minimum and average correlations with all human annotators except for themselves. All evaluations using MLLM are repeated three times, with the average value and standard deviation reported.
@@
Image Coherence MLLM 0.580.${}_{\text{0.023}}$ 0.440.${}_{\text{0.052}}$ 0.520.${}_{\text{0.036}}$
@@
Image Helpfulness MLLM 0.630.${}_{\text{0.032}}$ 0.450.${}_{\text{0.031}}$ 0.560.${}_{\text{0.021}}$
@@
Image Reference MLLM 0.670.${}_{\text{0.040}}$ 0.480.${}_{\text{0.031}}$ 0.590.${}_{\text{0.038}}$
@@
Since the reliability of text-modal metrics have been proven in RAG scenarios [5], we mainly focus on validating the reliability of our proposed MLLM-based multi-modal metrics: (1) Image Coherence; (2) Image Helpfulness; and (3) Image Reference. In practice, we randomly select 200 samples from the benchmark labeled by each metric, and three annotators independently scored these samples using the same scoring criteria as the model. Specifically, we calculate the Spearman correlation coefficients between two scorers, each can be human annotators or MLLMs. Table 1 reveals that MLLMs achieve comparable performance to inner-correlation of human annotators. GPT-4o is even better than human average on helpfulness metric (0.57 > 0.53). These results indicate that our proposed multi-modal evaluation metrics are good proxy for human annotators.
@@ 5.2 Overall Experimental Results
Reasoners Single OpenAI o3-mini 80.8 89.5 91.9 90.0 73.3 63.2 64.6 79.1 79.1
DeepSeek-R1 87.4 87.6 92.3 85.2 77.9 62.0 50.4 85.2 78.5
@@
We conduct comprehensive experiments on our benchmark with 12 representative models, including OpenAI o3-mini, DeepSeek-R1, GPT-4o, DeepSeek-V3, Step-1o, Llama-3.1 & Llama-3.2-Vision series, Qwen2.5 & Qwen2-VL series (see Appendix G for details). Due to the high evaluation cost, we sample 200 queries from the benchmark dataset by category to compare the performance of a wider range of models. Meanwhile, we also conduct experiments on the entire dataset containing 1,000 queries to further validate our findings. The results are shown in Table 2 and Table 9, and we can observe several findings.
@@
Table 2 indicates that the most capable open-source and closed-source LLMs for the M2RAG task are Qwen2.5-72B-Instruct and GPT-4o, using a multi-stage strategy with LLM backbones. Among MLLM backbones, GPT-4o and Step-1o outperform all open-source models. For example, Step-1o surpassing Qwen2-VL-72B-Instruct by a large margin (83.0 > 79.9). The reasoners OpenAI o3-mini and DeepSeek-R1 significantly outperform other models in the single-stage approach. Despite these strong performances, there remains significant room for improvement, as even the best-performing model falls short of the ideal score (83.8 < 100.0).
Reasoners Excel in Single-stage Generation
In single-stage settings, reasoner models (e.g., OpenAI o3-mini and DeepSeek-R1) significantly outperform standard LLMs, owing to their inherent chain-of-thought reasoning that guides both text and image selection. However, this advantage narrows under the multi-stage strategy: the explicit image-interleaving stages compensate for standard LLMs’ lack of built-in reasoning, allowing them to close the gap.
Abnormal Performance of Open-source MLLMs Reflect Pretraining Differences
The MLLMs with comparable capabilities, Llama-3.2-Vision and Qwen2-VL series, show great performance gap in the M2RAG task. The poor performance pf Llama-3.2-Vision is likely due to its pretraining on single image-text pairs [7], which limits its ability to handle multiple images. Qwen2-VL, on the other hand, has multi-image pretraining and thus supports multi-image reasoning and demonstrates more coherent output in multi-modal RAG tasks [32].
LLMs generally outperform open-source MLLMs of similar size
Due to the multi-image confusion phenomenon, current MLLMs struggle with reasoning over multiple images, aligning with the findings of previous works [22]. In contrast, LLMs receive detailed image descriptions and do not directly process raw visual input, avoiding this issue.
@@
Regarding modeling approaches, separate modeling [41] proves significantly weaker than joint modeling, with even the worst-performing joint approach (GPT-4o, 74.9) surpassing separate modeling (74.1), highlighting the importance of modeling multi-modal interactions.
@@ 5.3 Improvements from Fine-tuning
We further conduct the supervised fine-tuning to improve the 7B-8B LLMs and MLLMs capabilities in M2RAG task. Please refer to Appendix H for more fine-tuning implementation details. The evaluation results are illustrated in Table 2. It can be observed that our fine-tuned LLMs and MLLMs exhibit significant improvements on all multi-modal evaluation metrics and the overall score. For example, the average improvements on the overall score is 19.2%. Since the text-modal metrics are not utilized for constructing training dataset, our fine-tuned models underperforms the baselines on some textual evaluation metrics, particularly Fluency. Though the text-modal metrics are not our primary focus, it is still worth exploring whether the involving of these metrics in training data filtering can further improve the overall performance of open-source LLMs and MLLMs.
@@ 6.1 Does Topic Affect Performance?
Figure 3: Average overall score across 10 topics.
@@ 6.2 Is Description Quality Important?
Table 3: Results of ablation study on the quality of image descriptions and the inclusion of auxiliary images. Text Avg. and Multi-Modal Precision Avg. represent the average values of text-modal metrics and multi-modal precision metrics, respectively. Recall and Overall scores for w/o Aux. Image is not reported due to changes in the images within the knowledge base.
@@
As described in Section 4.3, each image is transcribed into a detailed textual description with its context for LLMs. Here we conduct ablation studies on the quality of image description. As illustrated in Table 3, we can conclude that (1) the exclusion of context in description generation leads to a significant decline in model performance under most metrics; (2) the absence of details in descriptions leads to a notable drop in Image Recall, highlighting the critical role of details in guiding image selection for LLMs. Overall, incorporating context and details in the image descriptions contribute to GPT-4o’s performance to a measurable extent.
@@ 6.3 Are Auxiliary Images Helpful?
As mentioned in Section 4.1.2, auxiliary images are collected to address cases where relevant images are absent from the retrieved web pages. As illustrated in last row of Table 3 (w/o Aux. Image), removing these auxiliary images leads to a significant decline in generative performance in multi-modal metrics. This observation reflects the challenges in effectively integrating auxiliary images that lack contextual information with the textual content. Overall, the inclusion of auxiliary images significantly enriches the visual content and enhances the quality of the generated document.
@@ Appendix A Limitations
The proposed evaluation metrics might fail to capture all potential aspects of M2RAG evaluation. For example, pairwise ranking is more appropriate for assessing overall performance than Likert scale in M2RAG task, which lacks a standard optimal answer. Besides, due to the requirement of using LLMs and MLLMs to evaluate the quality of long-text multi-modal content across 8 dimensions [^7] Please refer to Table 7 for more details about evaluation cost., this leads to huge evaluation costs in large-scale evaluation. To balance the affordability and reliability of experiments on our proposed benchmark, we introduce a quantity-first approach with 1,000 evaluation samples in this paper. In future work, we plan to distill and build efficient text-modal and multi-modal metrics to support larger-scale benchmark testing.
Appendix B Ethical Considerations
Most of the task inputs in our benchmark and training dataset are sourced from publicly available datasets, ensuring that they pose no harm to individuals or groups. Furthermore, the text generated by large language models (LLMs) and multi-modal language models (MLLMs) is carefully curated and processed by human annotators to safeguard privacy and confidentiality. No personally identifiable information (PII) is included. However, it is important to note that texts from the ELI5 dataset [6] and multi-modal documents retrieved via Google Custom Search may contain harmful content or hate speech. Despite these potential risks, it is crucial to disclose the full scope of this research, as materials from ELI5 and Google Custom Search have been extensively used in safety research within the community. All annotators are compensated fairly, with an hourly wage of approximately \$4.76 USD, which exceeds the average hourly wage of \$3.13 USD on Amazon Mechanical Turk [8].
Appendix C Query Collection
To simulate the real-world user query solving problem, we collect diverse and high-quality user queries from the ELI5 dataset [6] (The dataset is under BSD License). The ELI5 corpus is particularly suited for our task due to its comprehensive collection of long-form, open-ended questions that necessitate detailed, multi-sentence responses. The diversity of topics in ELI5 presents an opportunity to reflect the capabilities of language models in the real-world scenarios. Subsequently, we conduct two steps to collect diverse and high-quality queries that are suitable for using multi-modal information as responses: (1) Query Filtering; and (2) Query Classification.
@@ 1. Society & Culture
Example: What is the significance of the Day of the Dead celebration in Mexican culture? 2. Science & Mathematics: Questions in this category often deal with complex concepts that can be enhanced with diagrams, charts, or images. These visual aids can help in understanding scientific phenomena or mathematical theories. Example: How does photosynthesis work in plants? 3. Health: Health-related questions often benefit from the use of images like anatomical diagrams, charts, or photos for better understanding of medical conditions, treatments, or fitness exercises. Example: What are the symptoms and treatment options for carpal tunnel syndrome? 4. Education & Reference: This category includes academic topics and general knowledge where explanatory images, instructional graphics, and reference tables can provide clarity and enhance learning. Example: What are the main parts of a plant cell and their functions? 5. Computers & Internet: Questions about technology, software, and internet usage can be more effectively answered with screenshots, infographics, and step-by-step visual guides. Example: How do I set up a Wi-Fi network at home? 6. Sports: Sports-related queries can be explained better with visual demonstrations, play diagrams, athlete photos, and equipment images to illustrate techniques, rules, or historical moments. Example: What are the basic rules of soccer? 7. Business & Finance: Financial and business concepts often involve data, charts, and visualizations that can make complex information more digestible and insightful. Example: How does compound interest work and what impact can it have on investments? 8. Entertainment & Music: This category includes questions about movies, TV shows, music, and celebrities where images of album covers, scene snapshots, or charts of music theory can enrich the content. Example: What was the cultural impact of the TV show ‘Friends’?
@@ 11 Others 451
Appendix D Retrieval of Collected Elements
@@ 2. Text from a Web Page # Evaluation Guidelines
```markdown
{webpage_piece} ``` # Precautions
@@
Appendix E Dataset Statistics
E.1 Basic Statistics
The statistical information about our benchmark are shown in Table 5 and 6.
Table 5: The average values and standard deviations for element scores of the full benchmark dataset.
Item Avg. Score Std.
Text 1.5 2.3
Image 3.2 2.3
Aux. Image 3.8 2.6
Table 6: Numerical statistics of the full benchmark dataset.
Item Range Avg. Num Std.
Web Page per query 9.8 1.0
Image 12.0 16.7
Aux. Image 3.0 2.8
Text Element per web page 25.7 43.2
Image 1.2 4.9
E.2 Data Distribution
@@
Appendix F Details of Evaluation Metrics
F.1 Prompt Templates
@@
```markdown
{text} ``` # Scoring Criteria
@@
```markdown
{text} ``` # Scoring Criteria
@@
```markdown
{context_above} ``` # Image
@@
```markdown
{context_below} ``` # Scoring Criteria and Examples
@@
```markdown
{context_above} ``` # Image
@@
```markdown
{context_below} ``` # Scoring Criteria and Examples
@@
```markdown
{context_above} ``` # Image
@@
```markdown
{context_below} ``` # Scoring Criteria
@@ F.2 Evaluation Costs
Appendix G Experimental Setup
@@
We evaluate the capabilities of some advanced open-source and closed-source LLMs and MLLMs for M2RAG task. The involved models are listed in Table 8. Besides, we also implement a separate approach, i.e., the MuRAR method with GPT-4o as the backbone model [41].
Table 8: LLMs and MLLMs used for M2RAG task.
@@
For the generation process, we deployed open-source LLMs and MLLMs using vLLM [13] on Nvidia A100-SXM4-80GB GPUs. Models with 7B-11B parameters operate on a single GPU, while those exceeding 70B parameters are distributed across 4 GPUs using tensor parallelism. The context length is configured at 64k for the Llama series and 32k for the Qwen series. During generation, we use a top-$k$ selection method with $k$ set to 20 for filtering text elements from each webpage. The maximum number of auxiliary and web page specific images is restricted to 5, with a total input image cap at 10. Images are resized to 512 $\times$ 512 thumbnails for MLLMs processing. For evaluation, we implemented our customized metrics using the GPT-4o model to ensure the robust comprehensive assessment. Besides, the RAG metrics are implemented with RAGAS by using the GPT-4o-mini model.
Appendix H Implementation Details of Model Distillation
We fine-tune the open-source LLMs Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct and MLLM Qwen2-VL-7B-Instruct to serve as the distilled models. To ensure the fine-tuned model effectively captures the comprehensive information embedded in the training corpus, we set the context length to 32,768 tokens during fine-tuning, accommodating the majority of samples within the dataset. To optimize the computational efficiency and uphold the performance of the fine-tuned model, we employed Low-Rank Adaptation (LoRA) [10] with the rank of 128 and $\alpha$ of 256. Apart from that, we adopt various methods to accelerate training including ZeRO [24], Flash Attention 2 [4] and Liger Kernel [9]. The model training was conducted on 2 Nvidia A100-SXM4-80GB GPUs with a global batch size of 64 over 3 epochs, resulting in a total of 75 training steps. All models are fine-tuned with LLaMA-Factory framework [40].
Flu. Rel. CP. Faith. Coher. Help. Ref. Recall
Reasoners Single OpenAI o3-mini 83.8 90.7 86.7 86.5 76.1 65.6 70.2 85.2 80.6
DeepSeek-R1 86.5 87.7 87.1 77.8 78.2 62.9 53.3 88.8 77.8
Multi OpenAI o3-mini 82.6 85.6 90.9 81.6 72.4 63.2 81.4 100.0 82.2
DeepSeek-R1 83.0 85.4 79.5 71.9 77.6 69.9 81.8 99.9 81.1
LLMs Single GPT-4o 85.1 88.8 90.1 86.9 72.0 60.9 47.2 86.4 77.2
DeepSeek-V3 84.9 89.6 85.6 84.4 67.8 55.6 44.4 89.8 75.3
Multi GPT-4o 83.8 86.6 91.4 83.4 72.0 63.0 80.8 100.0 82.6
DeepSeek-V3 83.7 86.9 90.0 85.2 70.7 61.8 79.7 99.9 82.2
MLLMs Single GPT-4o 85.1 89.1 89.1 85.9 71.6 60.2 46.2 87.4 76.8
Step-1o 88.2 91.1 89.1 86.5 68.8 56.4 48.7 88.9 77.2
Multi GPT-4o 83.5 87.5 91.1 81.9 70.9 61.6 80.4 100.0 82.1
Step-1o 85.2 88.6 91.6 86.8 74.3 64.3 83.8 97.9 84.1
Fine-tuned Models
LLMs Single Llama-3.1-8B-Instruct 75.4 79.6 89.3 77.6 69.6 61.7 79.7 97.7 78.8
Qwen2.5-7B-Instruct 79.2 82.1 89.0 79.5 69.4 61.0 79.1 93.3 79.1
MLLM Single Qwen2-VL-7B-Instruct 62.2 78.2 87.7 68.4 67.0 59.3 75.7 95.0 74.2
Appendix J Case Study